# Learning Manipulation Skills through Robot Chain-of-Thought with Sparse Failure Guidance

### v1

@@ 5.2 Ablation Study
In this section, we conduct ablation studies by removing each component of our algorithm (CoT interpretation, human failure guidance, and success regularization). The results indicate that the most contributing factor to our algorithm’s performance is success regularization, followed by human failure guidance. Each component contributes to the overall performance of our algorithm. Additionally, we conduct ablation studies comparing GPT-4V success labels with ground truth success labels, demonstrating the robustness of our GPT-4V labeling. Furthermore, we explore different reward design mechanisms, showing that they can also achieve high performance within our algorithm. Finally, experiments with all baselines incorporating success regularization reveal that our algorithm’s superiority is not solely due to success regularization but also the reward design mechanism.
@@
Figure 4: Final policy performance for ablations and our algorithm across 3 random seeds. The results show that success regularization is the most significant contributing factor in our algorithm, followed by human failure guidance. Each component contributes to the overall performance of our algorithm.
Our algorithm incorporates several critical components: robot chain-of-thought processing via GPT-4V, sparse human failure guidance, and success experience regularization. To evaluate the significance of each component, we conducted a series of experiments, systematically removing each component to observe its impact on performance. A decline in performance upon the removal of any component demonstrates its essential role in our algorithm.
Figure 4 illustrates the success rates of the final policy performance of our algorithm compared to its ablated versions. The labels for these ablations are: w/oReg (without success experience regularization), w/oHG (without sparse human failure guidance), and w/oCoT (without robot chain-of-thought task decomposition).
The results confirm the critical importance of each component in our algorithm, with success experience regularization being particularly pivotal. Human failure guidance is the second most important component, motivating the agent to consider potential failure cases. Additionally, robot chain-of-thought processing helps the agent better understand the process of manipulation tasks, further contributing to the performance of our algorithm.
@@
Figure 6: Ablation study for our algorithm with CoT as one single prompt or multiple prompts. The results show that single prompt for reward mechanism also works well.
Our algorithm incorporates GPT-4V to gather successful trajectories for policy regularization. To validate the robustness of our GPT-4V labeling in collecting success experiences, we conduct experiments comparing GPT-4V labels and ground truth labels. If the performance using GPT-4V labels is similar to that using ground truth labels, it would confirm that GPT-4V is robust in collecting success experiences.
@@
Figure 6 displays the training success rate curves comparing GPT-4V and ground truth (GT) labeling methods for success experience collection. Our findings show that the GPT-4V labeling method achieves performance comparable to the ground truth across various tasks, demonstrating the robustness of our GPT-4V success labels.
Question 3: what is the empirical difference between our algorithm and the one with a single chain-of-thought interpreted positive prompt?
Our algorithm computes the NCE-rewards based on the maximum distance between the moving window video and multiple positive prompts. This measurement indicates how far the moving window video is from the language goal derived from robot chain-of-thought interpretations. Conversely, our ablation method computes the NCE-rewards based on the distance between the moving window video and a single interpreted positive prompt. The NCE-reward function is shown as:
$\text{reward}=\frac{\exp(z_{v}\cdot z_{l}^{+}/\tau)}{\exp(z_{v}\cdot z_{l}^{+}/\tau)+\sum_{z^{\prime}_{l}\in{\{z_{l}^{-}\}}}\exp(z_{v}\cdot z^{\prime}_{l}/\tau)},$ (5)
where $z_{l}^{+}$ indicates the single positive prompt from the GPT-4V interpretation. For example, the prompt for opening a door is "the robot hand first approaches the door’s handle, then grasps it, pulls it back, and finally opens the door."
Figure 6 shows the comparison between our algorithm with multiple positive prompts and the single positive prompt ablation. The empirical results demonstrate that our algorithm achieves comparable results with a single positive prompt. It indicates that our algorithm is robust to the setting of multiple or single positive prompts.
Figure 7: Ablation study for all baselines with success experience regularization. The results show that our algorithm performs well not only because of the success regularization component but also due to the reward design mechanism.
Question 4: what is the empirical difference between our algorithm and all baselines with success experience regularization?
Question 1 shows that success experience regularization component plays Question 1 shows that the success experience regularization component plays the most important role in our algorithm. To evaluate the significance of the reward design mechanism, we conduct experiments applying success regularization to all baselines. If the results show that our algorithm still outperforms other baselines with success regularization, then it proves that our algorithm’s superiority is due not only to the success regularization but also to the reward design mechanism.
Figure 7 shows the ablation results comparing our algorithm with other state-of-the-art baselines that include success experience regularization. The results demonstrate that our algorithm outperforms other baselines on almost all tasks. This indicates that our algorithm’s superiority is due not only to the success regularization component but also to the reward design mechanism.
@@ 6 Conclusion
In this study, we present a novel algorithm for robotic manipulation skill acquisition. We propose that pre-trained vision-language models can instruct robots on prior knowledge essential for manipulation tasks. Our algorithm simplifies complex tasks and acquires manipulation skills through a robot’s chain-of-thought process with sparse failure guidance, mirroring the developmental learning trajectory of a child. Specifically, our algorithm is designed to extract intricate temporal visual cues from the robot’s thought process, enabling a nuanced understanding of the task at hand. It also incorporates a learning strategy by describing potential failure cases through sparse failure guidance, which enhances its manipulation policies. Through success experience regularization, our algorithm strengthens its policies, thereby improving performance. Empirical results confirm the effectiveness of our approach, demonstrating a distinct advantage over established baseline models. This research sets a new standard for mastering manipulation skills in robots, and we hope it can be applied to more robotic manipulation tasks in the future.
7 Appendix
7.1 Environment Setup
We use MetaWorld-v2 as our test robotic environments. We choose 10 tasks as by [ROT] to evaluate our algorithm and baselines. They are drawer close, window close, door close, faucet close, button press, drawer open, window open, door open, faucet open and soccer. Among these tasks, opening objects such as drawers, windows, faucets, and doors is significantly more challenging than closing them.
To reduce the exploration space, we set the gripper a new position. With varying random seeds, the objects like drawer, window, door, faucet, button, and soccer would vary its positions. The position coordinate are provided in Table 1.
Table 1: Position coordinates for robot gripper as setting with MetaWorld-v2.
Task Coordinate
Drawer Close $0,0.6,0.2$
Drawer Open $0,0.8,0.05$
Window Close $0.25,0.80,0.16$
Window Open $-0.15,0.77,0.16$
Door Close $-0.5,0.6,0.2$
Door Open $0,0.6,0.2$
Faucet Close 0.,0.6,0.10.,0.6,0.10 . , 0.6 , 0.1
Faucet Open 0.,0.6,0.10.,0.6,0.10 . , 0.6 , 0.1
Button Press $0,0.6,0.2$
Soccer $0.6,0$ 0.6 , 0.06
Since the default camera view is not suitable for our algorithms. We re-set the camera for training all the experiments. The principle for choosing camera is that we set the object and gripper in the view. The camera coordinates are provided in Table 2.
Table 2: Position coordinates and angles for camera as setting with MetaWorld-v2.
Task Coordinate Angle
Door Close $-1.1,-0.4,0.6$
Other Tasks $0.55,0.3,0.5$ $3.9,2.3,0.6$
7.2 GPT-4V Success Labeling Details
Figure 8: The successful trajectories are gathered through GPT-4V success labeling. GPT-4V would label the success preference to the final observation of agent’s rollout trajectory. If it deemed successful, the trajectory would be recorded into a replay buffer.
Our algorithm incorporates GPT-4V to gather successful trajectories for policy regularization. The process for GPT-4V analysis is outlined in Figure 8. Initially, we present GPT-4V with a scenario indicating a successful state (e.g., the door is open). We then provide a prompt: "Consider the following image; this is a simulation environment. In this image, the door of the box is open. We define it when the door is open, the output is 1. Otherwise, the output is 0." Subsequently, we present the final observation to GPT-4V and prompt, "What is the output of this image?" We collect the trajectory as successful when GPT-4V’s response is 1.
To reduce the query times with GPT-4V, we distill a classifier from GPT-4V labeling data. First, we use GPT-4V to label 300 pictures with diverse final observations. Next, we fine-tune a ResNet-18 classifier based on this dataset. Before training, we use some augmentation techniques such as random crop resize to augment the dataset. Finally, we use this classifier to classify the final observation as successful or not for instance.
7.3 Task Robot Chain-of-Thought Prompt
We use GPT-4V to interpret task instructions such as "open the door." With GPT-4V, this instruction is broken down into a series of sub-goals $z_{l}^{+}$ through a CoT process, such as "the robot hand approaches the door’s handle," "the robot hand grasps the door’s handle," "the robot hand pulls the door’s handle back," and "the robot hand opens the door." Consequently, we leverage the fine-grained task descriptions from GPT-4V interpretation to enhance the agent’s understanding of the tasks. The task robot CoT prompts are provided in Table 3.
Table 3: Prompts from robot chain-of-thought for task interpretation.
Task Robot Chain-of-Thought Prompts
Drawer Close the robot hand pushes the drawer;
the drawer is closed.
Drawer Open the robot hand moves to the top of the drawer’s handle;
the robot hand moves inside the drawer’s handle;
the robot hand pulls back the drawer’s handle;
the robot hand opens the drawer.
Window Close the robot hand moves to the left of the window’s handle;
the robot hand pushes the handle towards right;
the robot hand closes the window.
Window Open the robot hand moves to the right of the window’s handle;
the robot hand pushes the handle towards left;
the robot hand opens the window.
Door Close the robot hand pushes the door close;
the door is closed.
Door Open the robot hand is at the front of the door;
the robot hand moves to the left of the door’s handle;
the robot hand pulls back the door’s handle;
the robot hand opens the door
Faucet Close the robot hand moves to the left of the faucet;
the robot hand pulls the faucet towards right;
the faucet is closed.
Faucet Open the robot hand moves to the right of the faucet;
the robot hand pulls the faucet towards left;
the faucet is open.
Button Press the robot hand moves towards the red button;
the robot hand pushes the button;
the button is pressed.
Soccer the robot hand moves towards soccer;
the robot hand pushes soccer into the net;
the soccer is in the net.
7.4 Sparse Human Failure Guidance
Robots may encounter potential failures during learning. Offline human failure guidance helps the agent consider potential failure cases, improving the policy for robustly learning manipulation skills.
$z_{l}^{-}$ is a set of offline potential failure case descriptions consisting of 2 to 4 prompts. To enable the agent to recognize and adjust to potential failures, we provide sparse human failure guidance before training the agent. The sparse human failure guidance are provided in Table 4.
Table 4: Sparse human failure guidance.
Task Human Failure Guidance
Drawer Close the robot hand is moving away from the drawer;
the robot hand is at the left of the drawer;
the robot hand is at the right of the drawer.
Drawer Open the robot hand is at the right of the drawer’s handle;
the robot hand is at the left of the drawer’s handle;
the robot hand is under the drawer’s handle.
Window Close the robot hand is at the top of the window’s handle;
the robot hand is at the right of the window’s handle;
the robot hand is under the window’s handle.
Window Open the robot hand is at the top of the window’s handle;
the robot hand is at the left of the window’s handle;
the robot hand is under the window’s handle.
Door Close the robot hand is moving away from the door;
the robot hand is at the left of the door.
Door Open the robot hand is at the right of the door’s handle;
the robot hand is at the top of the door’s handle;
the robot hand is under the door’s handle;
the robot hand is at the left of the box.
Faucet Close the robot hand is to the right of the faucet;
the robot hand is at the top of the faucet.
Faucet Open the robot hand is to the left of the faucet;
the robot hand is at the top of the faucet.
Button Press the robot hand moves away from the red button;
the robot hand is at the left of the red button;
the robot hand is at the right of the red button.
Soccer the robot hand is moving away from soccer;
the robot hand is at the top of soccer.

### v2

@@ 5.2 Ablation Study
In this section, we conduct ablation studies by removing each component of our algorithm (CoT interpretation, human failure guidance, and self imitation). The results indicate that the most contributing factor to our algorithm’s performance is VLM-based self imitation procedure, followed by human failure guidance. Each component contributes to the overall performance of our algorithm. Additionally, we conduct ablation studies comparing GPT-4V success labels with ground truth success labels, demonstrating the robustness of our GPT-4V labeling. Finally, experiments with all baselines incorporating VLM-based self imitation procedure reveal that our algorithm’s superiority is not solely due to self imitation but also the reward design mechanism.
@@
Figure 4: Final policy performance for ablations and our algorithm. The results show that VLM-based self imitation procedure is the most significant contributing factor in our algorithm, followed by human failure guidance.
Our algorithm incorporates several critical components: robot chain-of-thought, sparse human failure guidance, and VLM-based self imitation. To evaluate the significance of each component, we conducted a series of experiments, systematically removing each component to observe its impact on performance. A decline in performance upon the removal of any component demonstrates its essential role in our algorithm.
Figure 4 illustrates the success rates of the final policy performance of our algorithm compared to its ablated versions. The labels for these ablations are: w/oReg (without success experience regularization, i.e., self imitation), w/oHG (without sparse human failure guidance), and w/oCoT (without robot chain-of-thought task decomposition).
The results confirm the critical importance of each component in our algorithm, with VLM-based self imitation being particularly pivotal. Human failure guidance is the second most important component, motivating the agent to consider potential failure cases. Additionally, robot chain-of-thought processing helps the agent better understand the process of manipulation tasks, further contributing to the performance of our algorithm.
@@
Figure 6: Ablation study for all baselines with success experience regularization. The results show that our algorithm performs well not only because of the VLM-based self imitation component but also due to the reward design mechanism.
Our algorithm RoboCoT incorporates GPT-4V to gather successful trajectories for VLM-based self imitation procedure. To validate the robustness of our GPT-4V labeling in collecting success experiences, we conduct experiments comparing GPT-4V labels and ground truth labels. If the performance using GPT-4V labels is similar to that using ground truth labels, it would confirm that GPT-4V is robust in collecting success experiences.
@@
Figure 6 displays the training success rate curves comparing GPT-4V and ground truth (GT) labels for success experience collection. Our findings show that the GPT-4V labels achieves comparable performance to the ground truth across various tasks, demonstrating the robustness of our GPT-4V labels.
Question 3: what is the empirical difference between our algorithm and all baselines with VLM-based self imitation?
Question 1 shows that VLM-based self imitation component plays the most important role in our algorithm. To evaluate the significance of the reward design mechanism, we conduct experiments applying VLM-based self imitation procedure to all baselines. If the results show that our algorithm still outperforms other baselines with self imitation procedure, then it proves that our algorithm’s superiority is due not only to the VLM-based self imitation procedure but also to the reward design mechanism.
Figure 6 shows the ablation results comparing our algorithm with other state-of-the-art baselines that include VLM-based self imitation procedure. The results demonstrate that our algorithm outperforms other baselines on almost all tasks. This indicates that our algorithm’s superiority is due not only to the VLM-based self imitation component but also to the reward design mechanism.
@@ 6 Conclusion
In this study, we present a novel algorithm RoboCoT, which defines rewards for skill acquisition. Our algorithm RoboCoT simplifies complex tasks through a robot’s chain-of-thought process with sparse failure guidance. It yields a more fine-grained reward guidance. Through VLM-based self imitation, our algorithm strengthens its policies, thereby improving performance. Empirical results confirm the effectiveness of our approach, demonstrating a distinct advantage over established baseline models. This research sets a new standard for mastering manipulation skills in robots, and we hope it can be applied to more robotic tasks in the future.
7 Limitation
Our proposed method, RoboCoT, effectively masters a wide range of manipulation tasks through a robot chain-of-thought with sparse failure guidance. It provides a fine-grained reward guidance for the robot agent, facilitating the training of various automation tools. However, these tasks primarily focus on aligning robot movements with detailed language descriptions. RoboCoT does not yield promising results in tasks that emphasize observational preferences, such as soccer.